OpenStudy (anonymous):

how do i find the y-intercept of y=12x+120?

OpenStudy (anonymous):

Set x equal to zero and solve for y.

myininaya (myininaya):

since it is y=mx+b form b is your y intercept

OpenStudy (anonymous):

wait so whos right?

myininaya (myininaya):

both

myininaya (myininaya):

either way is fine for lines

OpenStudy (anonymous):

for a linear equation thought?

OpenStudy (anonymous):

your y intercept is 120

myininaya (myininaya):

this is a linear equation

OpenStudy (anonymous):

thanksss

myininaya (myininaya):

anything in this form y=mx+b or anything that can be written this way is a line

OpenStudy (anonymous):

for a linear equation thought?

OpenStudy (anonymous):

y=mx+b b is your y intercept so your y intercept is 120

myininaya (myininaya):

lol this is a line or a linear equation I don't know what you mean xoxo